mathslover (mathslover):

\(a^3 + b^3 + c^3 = 0 \) then find : \((a+b+c)_{(\textbf{min.})}\)

mathslover (mathslover):

\(a^3 + b^3 + c^3 = 0 \) then find : \((a+b+c)_{(\textbf{min.})}\) Justify your each step also.

mathslover (mathslover):

@uri sorry I edited the question, so you have to refresh your page.

mathslover (mathslover):

Remember I will ask for the proof of each and every step.

mathslover (mathslover):

\(\textbf{a,b and c} \in \mathbb{R}\)

Parth (parthkohli):

\[a^3 + b^3 + c^3 = (a + b + c) (a^2 + b^2 + c^2 - ab - bc - ca) + 3abc\]So \(abc = 0\) and \(a + b + c = 0\).

Parth (parthkohli):

Hmm... no

OpenStudy (anonymous):

good start point parth :)

Parth (parthkohli):

@mukushla Is that right?

OpenStudy (anonymous):

nope :)

Parth (parthkohli):

I think it should be \(abc = 0\) OR \(a + b + c = 0\).

OpenStudy (anonymous):

but the identity is good

mathslover (mathslover):

:)

Parth (parthkohli):

OR \(a^2 + b^2 + c^2 -ab-bc-ca =0\).

Parth (parthkohli):

I wonder if that even helps.

Parth (parthkohli):

\[a^3 + b^3 + c^3 = (a + b)(a^2 - ab + b^2) + c^3\]How about this one, maybe?

OpenStudy (anonymous):

Nope, why can you say abc=0 or a+b+c=0? It's not true from that identity.

Parth (parthkohli):

It's actually something like this: either \(abc = 0\) and \(a + b + c = 0\) or \(abc = 0\) and \(a^2+b^2+c^2−ab−bc−ca=0\). or \(abc = 0\) and \(a + b + c = 0\) and \(a^2+b^2+c^2−ab−bc−ca=0\).

OpenStudy (anonymous):

abc can be positive and (a+b+c) can be negative or vice versa, a^3+b^3+c^3 can still be zero.

Parth (parthkohli):

hmm

Parth (parthkohli):

You're right.

mathslover (mathslover):

;)

Parth (parthkohli):

Example?

OpenStudy (anonymous):

Have you learnt Lagrange multipliers @mathslover ?

Parth (parthkohli):

:-|||

Parth (parthkohli):

@drawar Are you Indian?

OpenStudy (anonymous):

Nope, why do you come up with that question?

mathslover (mathslover):

No but you can put your opinion here.

Parth (parthkohli):

Nothing, your name seemed Indian.

OpenStudy (anonymous):

Haha, just a random name, it has nothing to do with my real name.

OpenStudy (anonymous):

it's a method of finding the maximum/minimum values of functions subject to a number of constraints

mathslover (mathslover):

oh, yeah I remember I saw its name somewhere and turned over the page :P

OpenStudy (uri):

Hmmm

mathslover (mathslover):

Any easier method you have?

OpenStudy (shubhamsrg):

(a+b+c)/3 >= (abc)^1/3 and (a^3 + b^3 + c^3)/3 >= abc

OpenStudy (shubhamsrg):

is 0 the ans ?

mathslover (mathslover):

Yes ... 0 is the answer shubham .

OpenStudy (shubhamsrg):

well is my my method satisfactory ?

mathslover (mathslover):

It is easy that a+b+c will have 0 as the min. value but how will u prove that?

OpenStudy (shubhamsrg):

I used AM>=GM

mathslover (mathslover):

The expressions are right but how can u say that it is 0?

OpenStudy (shubhamsrg):

(a+b+c)/3 >= (abc)^1/3 and (a^3 + b^3 + c^3)/3 >= abc In second eqn, 0 >= abc so max value of abc is 0 and a+b+c>= 3 (abc)^1/3 when abc is max, a+b+c is min and hence the answer.

mathslover (mathslover):

\(\color{blue}{\textbf{EXCELLENT!}} \quad \ddot \smile\)

mathslover (mathslover):

Great work. @shubhamsrg

OpenStudy (shubhamsrg):

Wel, glad I could contribute

OpenStudy (shubhamsrg):

well*

mathslover (mathslover):

:)

mathslover (mathslover):

You deserve medals bro. Good work.